# Min-K%++: Improved Baseline for Detecting Pre-Training Data from Large Language Models

### v2

@@ Abstract
The problem of pre-training data detection for large language models (LLMs) has received growing attention due to its implications in critical issues like copyright violation and test data contamination. A common intuition for this problem is to identify training data by checking if the input comes from a mode of the LLM’s distribution. However, existing approaches, including the state-of-the-art Min-K%, often use zeroth-order signals for detection, which are less robust in determining local maxima than second-order statistics. In this work, we propose a novel methodology Min-K%++ for pre-training data detection that measures how sharply peaked the likelihood is around the input, a measurement analogous to the curvature of continuous distribution. Our method is theoretically motivated by the observation that maximum likelihood training implicitly optimizes the trace of the Hessian matrix of likelihood through score matching. Empirically, the proposed method achieves new SOTA performance across multiple settings. On the WikiMIA benchmark, Min-K%++ outperforms the runner-up by 6.2% to 10.5% in detection AUROC averaged over five models. On the more challenging MIMIR benchmark, it consistently improves upon reference-free methods while performing on par with reference-based method that requires an extra reference model.
@@ 1 Introduction
For these reasons, there has been growing interest in effective pre-training data detection strategies. Pre-training data detection can be considered a special case of Membership Inference Attack (MIA) [29]: the goal is to infer whether a given input has been used for training a target LLM (see Figure 1 left for illustration). Due to characteristics of pre-training corpora and training characteristics of LLMs [28, 10], this problem has been shown to be much more challenging than conventional MIA settings (see Section 5 for details).
In this work, we propose a novel methodology, Min-K%++, for pre-training data detection. The motivation lies in that existing methods (e.g., the state-of-the-art, Min-K[28]) often measure zeroth-order signals which could be less robust than second-order statistics in detecting training data (Section 3.1). Then, by connecting maximum likelihood training with score matching [17], we uncover that maximum likelihood training implicitly minimizes the trace of the Hessian matrix of likelihood over the training data (Section 3.2). This theoretical observation pinpoints a concrete second-order indicator for training data detection under a continuous distribution. Finally, we translate our insights into a practical and effective scoring mechanism that works for the discrete (categorical) distribution modeled by LLMs (Section 3.3). The core idea is to examine how sharply the likelihood is around the input, leading to an analogous measurement to the curvature of continuous distribution.
@@
Empirically, we perform extensive experiments to verify the proposed method on two established benchmarks, WikiMIA [28] and MIMIR [10]. Our experiments consider up to 5 families of models, including LLaMA [31], Pythia [2], and the new architecture Mamba [15]. Figure 1 right showcases the performance comparison between our method and previous ones. Notably, Min-K%++ on average leads to 6.2% to 10.5% absolute increases over the runner-up Min-Kon WikiMIA. On the more challenging MIMIR benchmark which consists of 7 subsets from various domains of the Pile dataset [13], our method still consistently outperforms Min-Kand is on par with the reference-based method [4]. Importantly, unlike reference-based methods which require another LLM to calibrate the likelihood, Min-K%++ is applied to a standalone target LLM. We also introduce and study an online detection setting that simulates “detect-while-generating”; the proposed Min-K%++ again performs the best. Lastly, Min-K%++ can also be interpreted from a calibration perspective: with ablation study we show that both calibration factors in the formulation contribute to the overall high performance.
@@ 2 Background
In this section, we first cover the problem statement of pre-training data detection defined by prior works [29, 28, 10]. Then, we go over the formulation of Min-K[28] as a representative method (due to its superior performance) to give a sense of how existing approaches handle the task.
@@
Min-Kmethod [28]. Recall that an auto-regressive LLM models the (conditional) categorical distribution of the next token given a prefix (i.e., all preceding tokens). Concretely, given a text input $x$ with $T$ tokens $(x_{1},x_{2},...,x_{T})$, we can obtain the probability $p(x_{t}|x_{<t})$ for each token $x_{t}$ given the prefix $x_{<t}$. [^1] For simplicity, from now on we use $(x_{<t},x_{t})$ to represent the token sequence $x_{1},x_{2},...,x_{t}$. Basing itself on a simple intuition that training texts are less likely to include low-probability “outlier” tokens, Min-Kdirectly measures the log probability of each token as the score with Equation 2. To aggregate the score for the full sentence, Equation 3 picks the $k$of tokens that have the lowest log probability and computes the average log probability over the selected tokens, with $k\in(0,100]$ being a hyperparameter. Formally,
@@
[…]
3 Min-K%++
In this section we introduce our methodology, called Min-K%++. We start by analyzing shortcomings of existing methods, where we point out that the zeroth-order signals measured by them may not be robust for detecting training data; instead, we argue that second-order statistic serves as a more reliable indicator. Then, by connecting the maximum likelihood training with score matching [17], we present a theoretical insight showing that the trace of the Hessian matrix of likelihood (which is second-order) is implicitly optimized on the training data. Finally, translating our insights into a practical solution, we formulate the proposed Min-K%++ which works for the discrete distribution of LLMs and measures an analogy of the curvature of continuous distribution.
3.1 Motivation: zeroth-order v.s. second-order
Intuitively, Min-Kand many existing methods try to identify training data by determining if the input comes from one of the modes of the probability distribution characterised by the model [28, 4]. However, they typically evaluate zeroth-order signals which are less robust than second-order statistics in telling whether the data forms a local maximum, a fact that we know from the derivative test [33]. Taking Min-Kas a representative example, it measures the absolute token probability as in Equation 2 which is only good for finding global maximum rather than local maximum. Later we will present a specific example to showcase such insufficiency.
3.2 Theoretical foundation: Hessian trace is optimized during maximum likelihood training
Pursuing along our motivation, we now wonder if there is a concrete second-order statistic that can help identify training data. Specifically, our exploration in this section features the key theoretical insight of this work, where we uncover that the trace of the Hessian matrix of likelihood (w.r.t. the input) is implicitly optimized through the course of maximum likelihood training. This observation lays solid foundation for later designing our methodology.
Our idea here is to ask an insightful question that, to our knowledge, previous works fail to touch on. That is, what effects or “footprints” do training samples leave to the model after training? To answer the question, first recall that (auto-regressive) LLMs are trained by maximizing the likelihood of each token $x_{t}$ conditioned on its prefix $x_{<t}$, i.e., $\log p(x_{t}|x_{<t})$. Meanwhile, for clarity, let us generalize from LLMs to the general maximum likelihood training for now.
Earlier, Koehler et al. [19] have proved that implicit score matching (ISM) [17] is a relaxation (within a multiplicative factor) of maximum likelihood estimation. Consequently, one can interpret maximum likelihood training loss through the lens of score matching and reformulate it with ISM [17] as
$\frac{1}{N}\sum_{\bm{x}}\biggl{[}\frac{1}{2}||\psi(\bm{x})||^{2}+\underbrace{\sum_{i=1}^{d}\frac{\partial\psi_{i}(\bm{x})}{\partial\bm{x}^{i}}}_{\text{Hessian trace}}\biggr{]},$ (4)
where $\psi(\bm{x})=\frac{\partial\log p(\bm{x})}{\partial\bm{x}}$ is the score function [17] for the $d$-dimensional input $\bm{x}$, $\psi_{i}(\bm{x})=\frac{\partial\log p(\bm{x})}{\partial\bm{x}^{i}}$, $\bm{x}^{i}$ is the element at the $i$-th dimension of $\bm{x}$, and $N$ is the number of training samples. Without ambiguity, we are omitting the model parameters in the formulation of $\psi$ for brevity.
Remark. Notice that the second term in the summation in Equation 4 is the trace of the Hessian matrix of log likelihood. Therefore, the key takeaway here is that (for continuous distributions) maximum likelihood training implicitly minimizes the Hessian trace of likelihood. More specifically, training data will tend to have smaller Hessian trace than non-training data, pinpointing a second-order signal that can provide accurate information for detecting training instances. This observation also in general applies to LLMs (as they are trained with maximum likelihood estimation), with a small caveat being that LLMs model a discrete (categorical) distribution instead of a continuous distribution. We address this caveat when formulating our method, which is discussed in the next section.
3.3 Formulating Min-K%++
Now that we have identified that the second-order Hessian trace is optimized over training data and thus encodes rich information for training data detection, we proceed to actually designing our scoring mechanism. Note, naïvely measuring the Hessian trace cannot work here because 1) the derivative is not well-defined for discrete inputs (i.e., tokens) of LLMs, and 2) even if it is somehow defined (e.g., defining the input as token embedding), computing gradients violates the grey-box access constraint.
Instead, our core idea is to measure how sharply peaked the likelihood is around an input, which takes inspiration from that Hessian matrix describes the curvature of a (continuous) function [35]. Concretely, we propose the following formulation for our Min-K%++:
[…]
@@ 3.3 Formulating Min-K%++
Here, $\mu_{x_{<t}}=\mathbb{E}_{z\sim p(\cdot|x_{<t})}[\log p(z|x_{<t})]$ is the expectation of the next token’s log probability over the vocabulary of the model given the prefix $x_{<t}$, and $\sigma_{x_{<t}}=\sqrt{\mathbb{E}_{z\sim p(\cdot|x_{<t})}[(\log p(z|x_{<t})-\mu_{x_{<t}})^{2}]}$ is the standard deviation. In practice, both terms can be computed analytically since the categorical distribution $p(\cdot|x_{<t})$ is encoded by the output logits of the model, which we have access to. Upon the token-wise score computed by Equation 5, we follow the same process as Min-Kto aggregate the sentence-wise score in Equation 6, i.e., we select the $k$of the tokens with the minimum score and compute the average score over them.
Interpretation 1: Measuring how sharply peaked the likelihood is. By doing this, our method puts up a measurement for discrete space that is analogous to the curvature of a continuous distribution. Specifically, the numerator of Equation 5 compares the likelihood $\log p(x_{t}|x_{<t})$, which corresponds to the ground-truth next token $x_{t}$, with the expected likelihood $\mu_{x_{<t}}$ computed over all the candidate tokens within the vocabulary. The larger $\log p(x_{t}|x_{<t})-\mu_{x_{<t}}$ is, the more probability mass gets assigned to the input token $x_{t}$ compared to other plausible tokens, meaning that $x_{t}$ tend to form a mode given $x_{<t}$ according to the model’s distribution (thus $(x_{<t},x_{t})$ is more likely to be training data).
@@
To see this perspective in a more straightforward way, let us consider a conceptual example depicted in Figure 2, where we show two instances of the distribution $p(\cdot|x_{<t})$ over a vocabulary of size 10. In the two cases (Figure 2 (a) and (b)), both $p(x_{t}|x_{<t})$ is 0.2, meaning that Min-Kwith this zeroth-order information will assign exactly the same score to the token $x_{t}$ and give the same judgement in both cases (e.g., that the sequence $(x_{<t},x_{t})$ was not seen in the training). However, inspecting them with the idea of our Min-K%++ reveals their distinct characteristics:
@@
  • In Figure 2 (a), one can notice that the distribution is actually peaked at $x_{t}$ (the bar with the hatch pattern ‘//’), indicating that the current sequence $(x_{<t},x_{t})$ forms a mode and is very likely to have appeared in the training.
@@
The above reasoning can be concisely reflected by comparing $\log p(x_{t}|x_{<t})$ and $\mu_{x_{<t}}$, which are marked by the two dashed lines in Figure 2. Accordingly, it can be seen that Figure 2 (a) and (b) exhibit positive and negative $\log p(x_{t}|x_{<t})-\mu_{x_{<t}}$, respectively, which successfully separate them apart. In addition to this high-level understanding, in Appendix A we also analytically show that the numerator shares similar form to an approximation of the Hessian trace, which further validates its design.
Interpretation 2: A calibration perspective. Reference-based methods [4] show superior performance as membership inference attacks by calibrating the (sentence-level) likelihood with certain references, such as the likelihood of the same input on a smaller LLM, the Zlib entropy, or the likelihood of lowercased text. In this regard, Min-K%++ can be seen as calibrating the token likelihood with two calibration factors, $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$. The unique advantage of our method lies in that instead of using some external references (e.g., an extra LLM), we leverage the statistics that can be readily computed with the target model only. The effect of $\mu_{x_{<t}}$ has been thoroughly discussed. The inclusion of $\sigma_{x_{<t}}$ is inspired by temperature scaling [16], a technique for calibrating the prediction confidence of neural networks. It scales the model output by a constant and has been shown to benefit tasks like Out-of-Distribution detection [20, 39]. Instead of using a constant, here we use a dynamic factor $\sigma_{x_{<t}}$ that is adaptive to each input and find that it leads to extra performance boosts. Later we will demonstrate that both factors contribute to the success of Min-K%++ through ablation study.
Figure 2: A conceptual example showcasing the idea of Min-K%++. (a) and (b) show two types of categorical distribution over the vocabulary. The bars with the hatch pattern (‘//’) correspond to the target token $x_{t}$. Min-K%++ measures how sharply peaked the likelihood is at the ground-truth next token $x_{t}$ by comparing $\log p(x_{t}|x_{<t})$ with $\mu_{x_{<t}}$.

### v3

@@ Abstract
The problem of pre-training data detection for large language models (LLMs) has received growing attention due to its implications in critical issues like copyright violation and test data contamination. Despite improved performance, existing methods (including the state-of-the-art, Min-K%) are mostly developed upon simple heuristics and lack solid, reasonable foundations. In this work, we propose a novel and theoretically motivated methodology for pre-training data detection, named Min-K%++. Specifically, we present a key insight that training samples tend to be local maxima of the modeled distribution along each input dimension through maximum likelihood training, which in turn allow us to insightfully translate the problem into identification of local maxima. Then, we design our method accordingly that works under the discrete distribution modeled by LLMs, whose core idea is to determine whether the input forms a mode or has relatively high probability under the conditional categorical distribution. Empirically, the proposed method achieves new SOTA performance across multiple settings. On the WikiMIA benchmark, Min-K%++ outperforms the runner-up by 6.2% to 10.5% in detection AUROC averaged over five models. On the more challenging MIMIR benchmark, it consistently improves upon reference-free methods while performing on par with reference-based method that requires an extra reference model.
@@ 1 Introduction
For these reasons, there has been growing interest in effective pre-training data detection strategies. Pre-training data detection can be considered a special case of Membership Inference Attack (MIA) [30]: the goal is to infer whether a given input has been used for training a target LLM (see Figure 1 left for illustration). Due to characteristics of pre-training corpora and training characteristics of LLMs [29, 10], this problem has been shown to be much more challenging than conventional MIA settings (see Section 5 for details). There are a few methods proposed recently, dedicated to this problem [4, 23, 29]. However, despite their improved performance, most existing methods are developed with simple heuristics, lacking solid and interpretable foundations.
In this work, we propose a novel and theoretically motivated methodology, Min-K%++, for pre-training data detection. Our exploration starts by asking a fundamental question of “what characteristics or footprints do training samples exhibit within the model”, where we answer it by revisiting the maximum likelihood training objective through the lens of score matching [18]. Importantly, we uncover that for continuous distributions with maximum likelihood training, training data points tend to be local maxima or be located near local maxima along input dimensions, a key insight that allows us for the first time to translate the training data detection problem into identification of local maxima. Then, upon theoretical insight, we develop a practical scoring mechanism that works for the discrete distribution modeled by LLMs. The core idea is to examine whether the input next token forms a mode or has relatively high probability under the conditional categorical distribution.
@@
Empirically, we perform extensive experiments to verify the proposed method on two established benchmarks with up to 5 families of models, including LLaMA [32], Pythia [2], and the new architecture Mamba [15]. Figure 1 right showcases the performance comparison between our method and previous ones. Notably, Min-K%++ on average leads to 6.2% to 10.5% absolute increases over the runner-up Min-K[29] on the WikiMIA benchmark [29]. On the more challenging MIMIR benchmark, our method still consistently outperforms other reference-free methods and is on par with the reference-based method [4]: unlike reference-based methods which require another LLM to calibrate the likelihood, Min-K%++ is applied to a standalone target LLM. We also introduce and study an online detection setting that simulates “detect-while-generating”; the proposed Min-K%++ again performs the best. Lastly, Min-K%++ can also be interpreted from a calibration perspective: with ablation study we show that both calibration factors in the formulation contribute to the overall high performance. We summarize our contributions as follows:
1) We establish that for maximum likelihood training, training instances tend to form local maxima or locate near local maxima along each input dimension (under continuous distributions), allowing us to turn training data detection into local maxima identification. 2) We develop a novel and sound methodology for LLM training data detection which functions by examining whether the input forms a mode or has relatively high probability under the conditional categorical distribution. 3) We demonstrate remarkable improvements over existing methods on established benchmarks.
@@ 2 Background
In this section, we first cover the problem statement of pre-training data detection defined by prior works [30, 29, 10]. Then we briefly introduce how LLMs work and how they are trained.
@@
(Auto-regressive) LLMs. LLMs are typically trained by maximum likelihood training that maximizes the probability of training token sequences [35]. In particular, auto-regressive LLMs decompose the probability of token sequence $(x_{1},x_{2},...,x_{t})$ with chain rule [37], i.e., $p(x_{1},x_{2},...,x_{t})=p(x_{t}|x_{1},x_{2},...x_{t-1})\cdot p(x_{1},x_{2},...x_{t-1})$. For brevity, throughout the paper we abbreviate the prefix of $x_{t}$ as $x_{<t}$. At inference stage, LLMs yield new tokens one by one according to its predicted conditional categorical distribution $p(\cdot|x_{<t})$ over the vocabulary.
3 Min-K%++
In this section we introduce our methodology, called Min-K%++. We start by discussing our grounding observation around general maximum likelihood training for continuous distribution, where we unveil that training samples tend to form local maxima or locate near local maxima along each input dimension of the distribution captured by the model. This insight allows us to introduce a unique and theoretically sound perspective for addressing the problem. Then, translating our insight into a practical solution, we formulate the proposed Min-K%++ which works for discrete inputs of LLMs by examining whether the input forms a mode or has a relatively high probability under the conditional categorical distribution of the predicted next token.
3.1 Motivation & Insight
To approach the pre-training data detection problem, unlike prior methods that often rely on heuristics [39, 5, 29], we choose to start our in-depth exploration by asking a fundamental question that existing works fail to touch to our knowledge. That is, what characteristics or “footprints” do training samples leave to the model after training? Since training data interacts with the model through training-time optimization, to answer the question, we revisit the maximum likelihood training objective of (auto-regressive) LLMs through the lens of score matching [18]. For the ease of discussion, let us generalize from LLMs, which model discrete distribution over tokens, to the general maximum likelihood training with continuous distribution for now.
Earlier, Koehler et al. [21] have proved that implicit score matching (ISM) objective [18] is a relaxation (within a multiplicative factor) of maximum likelihood estimation. Consequently, one can reformulate the maximum likelihood training loss with ISM as
$\frac{1}{N}\sum_{\bm{x}}\biggl{[}\frac{1}{2}||\psi(\bm{x})||^{2}+\underbrace{\sum_{i=1}^{d}\frac{\partial\psi_{i}(\bm{x})}{\partial\bm{x}_{i}}}_{\mathclap{\begin{subarray}{c}\text{the sum of the second-order}\\ \text{partial derivatives}\end{subarray}}}\biggr{]},$ (2)
where $\psi(\bm{x})=\frac{\partial\log p(\bm{x})}{\partial\bm{x}}$ is the score function [18] for a $d$-dimensional input $\bm{x}$, $\bm{x}_{i}$ is the element at the $i$-th dimension of $\bm{x}$, $\psi_{i}(\bm{x})=\frac{\partial\log p(\bm{x})}{\partial\bm{x}_{i}}$, and $N$ is the number of training samples. Without ambiguity, we are omitting the model parameters in the formulation of $\psi$ for brevity.
Remark. It can be seen from Equation 2 that maximum likelihood training implicitly minimizes both 1) the magnitude of first-order derivatives of likelihood $\log p(\bm{x})$ w.r.t. $\bm{x}$ and 2) the sum of the second-order partial derivatives of $\log p(\bm{x})$ w.r.t. each dimension of $\bm{x}$. As a result, we posit that the first-order $\frac{\partial\log p(\bm{x})}{\partial\bm{x}_{i}}$ will be close to 0 and the second-order $\frac{\partial^{2}\log p(\bm{x})}{\partial\bm{x}_{i}^{2}}$ will be minimized to be negative for training samples. Importantly, according to the univariate second-derivative test [34], the implication here is that when looking along each dimension of the input (corresponding to each $\bm{x}_{i}$), training data points tend to form local maxima or be located near local maxima of the likelihood. Given this key insight, we can translate the problem of detecting training data into identifying whether the input is a local maximum or locates near a local maximum (along input dimensions), providing a novel and theoretically grounded perspective for tackling this task. Note that although we focus on LLMs in this work, our discussion here is general and can be applied to other modalities and models, e.g., VAEs [19] and diffusion models [17, 28] for image modeling, which are trained by optimizing the evidence lower bound of likelihood [19, 17].
3.2 Formulating Min-K%++
From insights to an actual method. In the context of LLMs, the dimension of a token sequence corresponds to each token. Also, note that tokens are discrete rather than continuous and there is no notion of locality (local maxima) anymore. Therefore, to determine whether a token sequence $(x_{<t},x_{t})$ has appeared in training, we choose to examine whether it forms a mode or has a relatively high probability along the dimension that corresponds to $x_{t}$. Again, in our context, moving along the dimension of a token means replacing the current one with other candidate tokens within the vocabulary of the model. Our idea here closely reflects the insight in the continuous distribution case we have discussed above, where we proposed to detect training data by identifying if the input is a local maximum or locates near a local maximum.
Meanwhile, since the probability $p(x_{<t},x_{t})$ can be decomposed by $p(x_{<t},x_{t})=p(x_{t}|x_{<t})\cdot p(x_{<t})$, the problem further transforms into identifying if the next token $x_{t}$ forms a mode or has a relatively high probability under the conditional categorical distribution $p(\cdot|x_{<t})$. To this end, we propose the following formulation for our Min-K%++:
@@ 3.2 Formulating Min-K%++
Here, $\mu_{x_{<t}}=\mathbb{E}_{z\sim p(\cdot|x_{<t})}[\log p(z|x_{<t})]$ is the expectation of the next token’s log probability over the vocabulary of the model given the prefix $x_{<t}$, and $\sigma_{x_{<t}}=\sqrt{\mathbb{E}_{z\sim p(\cdot|x_{<t})}[(\log p(z|x_{<t})-\mu_{x_{<t}})^{2}]}$ is the standard deviation. In practice, both terms can be computed analytically since the categorical distribution $p(\cdot|x_{<t})$ is encoded by the output logits of the model, which we have access to. Upon the sequence-wise score computed by Equation 3, we adopt a similar strategy to Shi et al. [29] where we select the $k$of the token sequences with the minimum score and compute the average over them to obtain the final sentence-wise score, as described by Equation 4. This is because in practice LLM training minimizes the loss of all possible sequences in each training text/sentence. Therefore, by computing an aggregated score we expect to get a more robust estimation for the full sentence input. Our method is closely motivated by our theoretical insight. We provide further interpretations for more comprehensive understanding.
Interpretation 1: More robust identification of the mode. We have translated training data detection into determining whether the input next token forms a mode of the conditional distribution. Besides motivating our method, this unique scheme also allows us to understand why existing methods may be sub-optimal (which was unclear earlier since they were mostly developed upon heuristics).
Specifically, let us take the state-of-the-art Min-K[29] as an example, which follows a simple intuition that training texts are less likely to include low-probability “outlier” tokens and thus uses the token probability as the score. Within our proposed scheme, this may not identify the mode robustly or accurately as token probabilities can vary a lot across different inputs. For example, some training texts may be rare or inherently difficult to learn, thus having low token probabilities despite still being the mode. Our design, in contrast, can account for such cases since it explicitly compares the log likelihood of the target next token $\log p(x_{t}|x_{<t})$ and the expected log likelihood over all candidate tokens $\mu_{x_{<t}}$. Regardless of the absolute value of the token likelihood, the larger $\log p(x_{t}|x_{<t})-\mu_{x_{<t}}$ is, the more likely $x_{t}$ forms a mode or has a relatively high probability among the vocabulary.
@@
To see this more clearly, let us consider a conceptual example depicted in Figure 2, where we show the predicted next token probability distribution for two different input token sequences $(x_{<t},x_{t})$, supposing a vocabulary of size 10. In the two cases (Figure 2 (a) and (b)), both $p(x_{t}|x_{<t})$ is 0.2. Existing methods often measure exactly the next token probability [29] and thus will give the same judgement to both inputs (e.g., that the sequence $(x_{<t},x_{t})$ was not seen in the training). However, inspecting them with the idea of our Min-K%++ reveals their distinct characteristics:
@@
  • In Figure 2 (a), one can notice that the distribution is actually peaked at $x_{t}$ (the bar with the hatch pattern ‘//’), indicating that the current sequence $x_{t}$ forms a mode of $p(\cdot|x_{<t})$ and is very likely to have appeared in the training.
@@
The above reasoning can be concisely reflected by comparing $\log p(x_{t}|x_{<t})$ and $\mu_{x_{<t}}$, which are marked by the two dashed lines in Figure 2. Accordingly, it can be seen that Figure 2 (a) and (b) exhibit positive and negative $\log p(x_{t}|x_{<t})-\mu_{x_{<t}}$, respectively, which successfully separate them apart.
Interpretation 2: A calibration perspective. Reference-based membership inference attacks [4] show superior performance by calibrating the (sentence-level) likelihood with certain “references”, such as the likelihood of the same input on a smaller LLM, the Zlib entropy, or the likelihood of lowercased text. In this regard, Min-K%++ can be seen as a calibrating method, too, where it calibrates the next token log likelihood with two calibration factors, $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$. The unique advantage of our method lies in that instead of using some external references (e.g., an extra LLM), we leverage the statistics that can be readily computed with the target model only. The effect of $\mu_{x_{<t}}$ has been thoroughly discussed in the first interpretation. The inclusion of $\sigma_{x_{<t}}$ is inspired by temperature scaling [16], a technique for calibrating the prediction confidence of neural networks. It scales the model output by a constant and has been shown to benefit tasks like out-of-distribution detection [22, 41]. Instead of using a constant, here we use a dynamic factor $\sigma_{x_{<t}}$ that is adaptive to each input and find that it leads to extra performance boosts. Later we will demonstrate that both factors contribute to the success of Min-K%++ through ablation study.
Figure 2: A conceptual example showcasing the idea of Min-K%++. (a) and (b) show two next token distributions for two different input token sequences $(x_{<t},x_{t})$. The bars with the hatch pattern (‘//’) correspond to the target next token $x_{t}$. Existing methods often measure the exact next token probability of $x_{t}$ (hatched bars), which fails to separate these two inputs due to the same probability (0.2). By comparison, Min-K%++ determines if $x_{t}$ forms a mode or has relatively high probability by comparing $\log p(x_{t}|x_{<t})$ with $\mu_{x_{<t}}$; the score is the difference between the red and pink dashed lines. See text for detailed discussion.